# Comprehensive Machine Learning Benchmarking for Fringe Projection Profilometry with Photorealistic Synthetic Data

### v1

@@ 4.2 Network Architectures
Hformer [30]: Hybrid CNN-transformer with HRNet-W18 backbone for multi-scale features [18,36,72,144], transformer encoder-decoder with window-based attention (size 8), patch expansion upsampling. Dropout 0.5, Adam optimizer with RMSE loss.
@@
ResUNet [9]: UNet with residual blocks replacing convolutional blocks, four levels (960960 to 120120), identity skip connections for gradient flow. Dropout 0.5, RMSProp optimizer (, lr= with ReduceLROnPlateau).
4.3 Training and Evaluation
All networks except Pix2Pix use Adam optimizer (lr=, =(0.9,0.999), weight decay=), with learning rate reduction by 0.1 after 10 plateau epochs. Training continues for max 1000 epochs with early stopping after 50 non-improving epochs.
Loss function is RMSE over normalized depth values:
where is the number of valid pixels (excluding background). Training uses batch size 4 (UNet) or 1 (Hformer) on NVIDIA A100 GPUs with mixed precision.
Models are evaluated after denormalizing to metric units (mm):
Results are reported as aggregate statistics across all 30 test samples.
4.4 Experimental Results
4.4.1 Quantitative Comparison
Table 2 shows three models achieve nearly identical performance (Pix2Pix: 58.89 mm, Hformer: 59.92 mm, UNet: 60.26 mm RMSE) despite architectural differences, while ResUNet underperforms (76.55 mm, 30% worse). Mean errors represent 74-96% of the typical 80 mm depth range. High standard deviations (26.71-29.65 mm, 50% of mean) reveal geometry-dependent brittleness: simple objects yield 9-15 mm errors while complex shapes produce 100-120 mm errors.
Figure 2 shows error distributions and per-sample errors. The curves track closely for Pix2Pix, Hformer, and UNet, while ResUNet consistently exhibits higher errors. Certain test samples present extreme difficulty for all models (80-120 mm errors), while others yield relatively low errors (10-30 mm), indicating reconstruction accuracy is dominated by geometric factors rather than model-specific capabilities.
4.4.2 Qualitative Analysis
Figure 3 shows representative results for one test object. Models capture coarse shape and approximate depth ordering but fail on fine details with exception of UNet and ResUNet. Error concentrates at boundaries and discontinuities, suggesting networks learn semantic shape completion rather than geometric reconstruction from fringe deformation. Predictions for Hformer in particular resemble smooth, regularized versions of true geometry, as if networks project ambiguous observations onto learned shape manifolds. Meanwhile, Pix2Pix is unable to predict the rough overall geometry of the object, while every other model can.
4.5 Discussion
Results demonstrate that direct single-shot fringe-to-depth regression without explicit phase extraction yields limited reconstruction accuracy. Mean errors of 59-77 mm (74-96% of 80 mm depth range) indicate networks learn coarse shape approximations rather than accurate geometry.
First, single fringe images lack sufficient information for accurate reconstruction. The periodic nature of sinusoidal patterns creates fundamental ambiguities unresolved by training on diverse geometries. Traditional FPP uses temporal redundancy or spatial unwrapping; our experiments show networks instead learn to map fringe patterns to plausible depth distributions based on statistical regularities.
Second, high variance across test objects (standard deviations 26.71-29.65 mm, 50% of mean) reveals geometry-dependent brittleness. Simple objects yield 9-15 mm errors while complex shapes produce 100-120 mm errors, suggesting single-shot approaches may only work for constrained domains with limited geometric variation.
Third, network architecture plays surprisingly limited role. Pix2Pix, Hformer, and UNet achieve nearly identical performance (58.89-60.26 mm RMSE) despite substantial design differences: adversarial vs supervised, hybrid CNN-transformer vs pure CNN. Although in the case of Hformer and UNet the overall shape of the object is captured, Pix2Pix fails to do that, which can be attributed the requirement of larger training data for Pix2Pix [6]. This suggests architectural innovations provide minimal benefit for this problem without addressing the fundamental information deficit. ResUNet’s underperformance (76.55 mm) despite increased depth suggests overfitting on the limited 240-sample training set.
These findings motivate incorporating explicit phase information as intermediate representations. Rather than end-to-end fringe-to-depth regression, future work should train networks to refine phase unwrapping, denoise wrapped phase, or predict depth from coarse phase estimates, leveraging geometric structure while benefiting from learned priors.
@@ 5 Conclusion and Future Work
This paper presents the first comprehensive machine learning benchmarking framework for FPP, creating a large-scale synthetic dataset (15,600 images, 300 reconstructions, 50 objects) with perfect ground truth using VIRTUS-FPP. Our benchmarking reveals four architectures achieve similar performance (58-77 mm RMSE) with errors approaching 75-95% of the depth range, demonstrating fundamental limitations of single-shot fringe-to-depth mapping without explicit phase information.
The near-identical performance across diverse architectures indicates the information deficit, not model design, limits reconstruction quality. Networks learn semantic shape priors rather than accurate geometry. These results strongly motivate hybrid approaches combining traditional phase-based FPP with learned refinement.
Future directions include: (1) Phase-guided learning using wrapped/unwrapped phase maps, (2) Sim-to-real transfer via domain adaptation leveraging VIRTUS-FPP’s digital twin capability, (3) Multi-view fusion using 6 viewpoints per object, (4) Task reformulation for post-processing traditional reconstructions, (5) Dataset expansion to challenging materials and lighting with domain randomization [19], and (6) Uncertainty quantification through probabilistic deep learning.
By providing comprehensive synthetic data and standardized evaluation protocols, this work establishes a foundation for systematic, data-driven FPP research, enabling development of robust systems for manufacturing, biomedical imaging, and automated inspection.
@@ Acknowledgements.
- [6] (2025) A dual-aspect evaluation framework for architectural-like plan generation via pix2pix series algorithms. Frontiers of Architectural Research. Cited by: §4.5.
@@
- [9] (2025) Deep-learning-assisted single-shot 3d shape and color measurement using color fringe projection profilometry. Optical Review, pp. 1–12. Cited by: §1, §4.2.
@@
- [17] (2015) U-Net: convolutional networks for biomedical image segmentation. In Medical Image Computing and Computer-Assisted Intervention (MICCAI), pp. 234–241. Cited by: §4.2.
@@
- [21] (2021) Single-shot fringe projection profilometry based on deep learning and computer graphics. Optics Express 29 (6), pp. 8024–8040. Cited by: §1.
@@
- [29] (2020) Fringe projection profilometry by conducting deep learning from its digital twin. Optics Express 28 (24), pp. 36568–36583. Cited by: §1, §1, §3.2.
- [30] (2022) Hformer: hybrid convolutional neural network transformer network for fringe order prediction in phase unwrapping of fringe projection. Optical Engineering 61 (9), pp. 093107–093107. Cited by: §1, §4.2.

### v2

@@ 4.4 Architecture Comparison
Hformer [33]: Hybrid CNN-transformer with HRNet-W18 backbone for multi-scale features [18,36,72,144], transformer encoder-decoder with window-based attention (size 8), patch expansion upsampling.
@@
ResUNet [9]: UNet with residual blocks replacing convolutional blocks, four levels ( to ), identity skip connections for improved gradient flow.
Pix2Pix [10]: Conditional GAN with U-Net generator and PatchGAN discriminator, adapted from NVIDIA Pix2Pix-HD [26].
| Architecture | Overall | Overall | Object | Object | Background | Background |
| MAE | RMSE | MAE | RMSE | MAE | RMSE | |
| UNet | 3.31 | 9.85 | 14.54 | 17.88 | 2.01 | 8.44 |
| Hformer | 3.62 | 14.55 | 22.17 | 26.78 | 1.72 | 12.87 |
| ResUNet | 3.73 | 13.73 | 23.32 | 28.38 | 1.80 | 11.54 |
| Pix2Pix | 5.48 | 12.20 | 27.73 | 38.22 | 3.16 | 5.15 |
Results: Table 5 and Figure 9 summarize performance across architectures. UNet achieves the best object-only error (14.54 mm MAE, 17.88 mm RMSE), outperforming Hformer by 52%, ResUNet by 60%, and Pix2Pix by 91%. Figure 10 shows representative predictions for a container bottle object.
UNet’s superior performance under the optimized configuration suggests that the baseline encoder-decoder architecture with skip connections is well-suited for this task when combined with proper normalization and loss function design. The consistent [0,1] normalized space and hybrid loss allow UNet’s simple architecture to focus learning on shape reconstruction without the complexities of handling scale variation or architectural sophistication.
Hformer and ResUNet achieve similar mid-range performance (22-23 mm object MAE), despite their architectural differences. Hformer’s transformer-based attention mechanisms and ResUNet’s residual connections provide comparable benefits, but neither architecture overcomes the fundamental information deficit of single-shot reconstruction. The 52-60% degradation compared to UNet suggests that architectural complexity may actually hinder performance on this task, potentially due to overfitting on the limited 240-sample training set or difficulty optimizing more complex models in the constrained [0,1] normalized space.
Pix2Pix performs worst (27.73 mm object MAE, 38.22 mm object RMSE), revealing a fundamental misalignment between adversarial training and metrological objectives. Figure 10 shows a counterintuitive result: Pix2Pix produces visually compelling predictions with smooth surfaces and statistically correct depth distributions—the container bottle prediction exhibits a depth range [1561.1, 1776.7] mm that exactly matches ground truth—yet achieves nearly 2 worse error than UNet (14.54 mm MAE). This occurs because the adversarial discriminator optimizes for perceptual plausibility (“does this look like a depth map?”) rather than point-wise accuracy. The network learns to match statistical properties (min, max, mean depth) and produce realistic surface geometries, but introduces smooth systematic offsets of 20-40 mm that are visually imperceptible yet catastrophic for precision measurement. These offsets appear as uniform color shifts in visualizations rather than high-frequency noise, making predictions appear deceptively accurate to human observers while failing metrological evaluation. The severe background artifacts further confirm this: the network learned that “depth maps should have varying depth values” without the critical constraint that background regions must predict exactly zero depth.
Critically, even the best architecture (UNet at 14.54 mm MAE) remains far from the sub-millimeter accuracy of traditional FPP. For objects with an 80 mm depth range in our dataset, this represents 18% relative error. The modest 1.9 performance gap between best (UNet: 14.54 mm) and worst (Pix2Pix: 27.73 mm) architectures, combined with all models’ failure to achieve precision depth reconstruction, confirms our hypothesis: the fundamental limitation is information deficit, not model design. Single fringe images lack sufficient information for accurate depth recovery, and no amount of architectural engineering can overcome this constraint.
This finding parallels the loss function results (Section 4.3), where even optimal loss design provided only 10% improvement over baseline. Together, these results demonstrate that single-shot fringe-to-depth mapping without explicit phase information is fundamentally limited. Networks learn coarse shape priors and statistical regularities rather than accurate geometry, regardless of architectural sophistication.
@@ 5 Conclusion and Future Work
This paper presents the first comprehensive open-source machine learning benchmarking framework for fringe projection profilometry, establishing both optimal learning configurations and fundamental limitations for single-shot 3D reconstruction. Using a large-scale synthetic dataset (15,600 fringe images, 300 depth reconstructions, 50 objects) generated with VIRTUS-FPP’s physics-based rendering. We conducted a systematic three-phase ablation study that reveals critical insights for learning-based FPP.
Phase 1 - Data representation: Individual normalization, which decouples object shape learning from absolute scale, achieves 9.1 better reconstruction (16.20 mm vs 148.07 mm object MAE) than raw depth and 5.1 better than global normalization. Unexpectedly, removing background fringe patterns degrades performance 2.8-7.3 across all normalizations, demonstrating that background fringes provide essential spatial phase reference rather than acting as noise. This finding challenges conventional assumptions about input preprocessing for FPP learning.
Phase 2 - Loss function optimization: Among six L1/L2-based losses, Hybrid L1 with =0.7 achieves best performance (14.54 mm object MAE, 10% improvement over baseline). Critically, masked losses that focus exclusively on object pixels catastrophically fail (122-147 mm overall MAE) due to scale drift, paralleling the background ablation finding: both removing background fringes physically and masking them computationally eliminate essential spatial constraints, implying background information is signal, not noise, for single-shot FPP.
Phase 3 - Architecture comparison: Using the obtained optimal configuration (individual normalization + Hybrid L1 =0.7), UNet achieves best performance (14.54 mm object MAE, 17.88 mm object RMSE), outperforming Hformer by 52%, ResUNet by 60%, and Pix2Pix by 91%. The modest 1.9 gap between best (14.54 mm) and worst (27.73 mm) architectures, combined with all models’ failure to achieve sub-millimeter accuracy, confirms that information deficit limits reconstruction quality. Pix2Pix’s counterintuitive failure reveals that its predictions exhibit statistically correct depth distributions (e.g., exact range match [1561-1777 mm]) and smooth, realistic surfaces, yet achieve worst metric accuracy. This reveals fundamental misalignment between adversarial training’s perceptual plausibility objective and metrology’s point-wise accuracy requirement, explaining why modern depth estimation do not use GANs.
Even the best model’s 14.54 mm error represents 18% of objects in our dataset with 80 mm depth range, far from traditional FPP’s sub-millimeter capabilities. This demonstrates that single fringe images fundamentally lack information for accurate depth recovery without explicit phase: the periodic nature of sinusoidal patterns creates inherent ambiguity that cannot be resolved through learned shape priors alone. Networks achieve coarse semantic reconstruction but fail on precise geometry.
These findings strongly motivate hybrid approaches combining traditional phase-based FPP with learned refinement. Future directions include: (1) Phase-guided learning using wrapped/unwrapped phase maps as network input or intermediate representations, (2) Multi-view fusion leveraging 6 viewpoints per object in our dataset, (3) Sim-to-real transfer via domain adaptation using VIRTUS-FPP’s digital twin capability, (4) Task reformulation for post-processing traditional reconstructions (denoising, hole-filling, outlier removal) rather than end-to-end depth prediction, (5) Dataset expansion to challenging materials (specular, translucent) and lighting conditions with domain randomization [21], and (6) Uncertainty quantification through probabilistic deep learning to identify when predictions are unreliable.
By providing a comprehensive open-source synthetic data, standardized evaluation protocols separating object from background errors, and systematic evidence of fundamental limitations, this work establishes a foundation for data-driven FPP research. Our findings suggest that rather than pursuing increasingly complex architectures for direct fringe-to-depth mapping, the field should focus on integrating learned components with physics-based reconstruction pipelines to leverage both traditional FPP’s geometric precision and neural networks’ robustness to noise and ambiguity.
@@ Acknowledgements.
- [5] (2021) Calibration of fringe projection profilometry: a comparative review. Optics and lasers in engineering 143, pp. 106622. Cited by: §1, §4.2.
@@
- [9] (2025) Deep-learning-assisted single-shot 3d shape and color measurement using color fringe projection profilometry. Optical Review, pp. 1–12. Cited by: §1, §1, §4.2, §4.2, §4.4.
@@
- [13] (2025) Deep-learning-enabled single-shot fringe projection profilometry based on inner shifting-phase encoding. Optics Express 33 (23), pp. 49530–49550. Cited by: §1.
@@
- [15] (2020) Single-shot 3d shape reconstruction using structured light and deep convolutional neural networks. Sensors 20 (13), pp. 3718. Cited by: §1, item 2.
@@
- [19] (2015) U-Net: convolutional networks for biomedical image segmentation. In Medical Image Computing and Computer-Assisted Intervention (MICCAI), pp. 234–241. Cited by: §4.2, §4.4.
@@
- [23] (2021) Single-shot fringe projection profilometry based on deep learning and computer graphics. Optics Express 29 (6), pp. 8024–8040. Cited by: §1, §1.
@@
- [25] (2025) End-to-end single-shot composite color fpp network for multiple separated objects reconstruction. Measurement 246, pp. 116697. Cited by: §1.
@@
- [32] (2020) Fringe projection profilometry by conducting deep learning from its digital twin. Optics Express 28 (24), pp. 36568–36583. Cited by: §1, §1, §3.2, §4.2.
- [33] (2022) Hformer: hybrid convolutional neural network transformer network for fringe order prediction in phase unwrapping of fringe projection. Optical Engineering 61 (9), pp. 093107–093107. Cited by: §1, §1, §4.4.